Circular No. 8948

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V5579 SAGITTARII R. W. Russell, R. J. Rudy, and D. K. Lynch, The Aerospace Corpora-

tion; C. E. Woodward, University of Minnesota; and H. Marion, University of Texas and Texas State University, report observations of V5579 Sgr (cf. IAUC 8937, 8940) on May 9 UT at the Infrared Telescope Facility (+ SPEX; range 0.8–5.2 μm). D. Griep, IRTF, helped to acquire the data. The nova was at a very early stage of its spectral development, showing lines of O I, N I, and Ca II, and exceptionally strong lines of C I. Lines widths were ≈ 1600 km/s at FWHM. Fe II features were weak, but — despite that fact — this is almost certainly an “Fe II”-type of nova. Lines of neutral helium had not yet formed. The strongest lines in the infrared spectrum are the O I lines that are fluourescently excited by Lyβ. These lines indicate a reddening of E(B–V) = 1.2, some of which may be local to the nova, which has already formed dust. In fact, the infrared continuum is dominated by thermal emission from dust at a single temperature of 1370 K.

NOVA OPHIUCHI 2008 K. Ayani and N. Murakami, Bisei Astronomical Observatory (BAO),

write that low-resolution spectra (range 400–800 nm; resolution 0.5 nm at Hα) of the apparent nova announced on IAUC 8947, obtained with the BAO 1.01-m telescope on May 26.7 UT, confirms that the object is a nova. They show Hα emission (FWHM ∼ 600 km/s) with a P-Cyg profile, its absorption minimum being blueshifted by 730 km/s with respect to the emission peak. Hβ and Fe II (multiplet 42) lines also have P-Cyg profiles, although the absorption component dominates on lines of shorter wavelength.

Further to IAUC 8947, S. Nakano reports that co-discoverers K. Nishiyama and F. Kabashima obtained additional unfiltered CCD frames with a 0.40-m reflector on May 26.562 UT, yielding mag 9.8 for the nova and position α = 17^{h}39^{m}50^{s}.94, δ = −23^{o}50′01′′.0 (equinox 2000.0). Nakano also forwards an unfiltered CCD observation by K. Kadota (Ageo, Japan, 0.25m reflector) on May 26.609 that yielded mag 9.9 and a measured position identical to that above.

Visual magnitude estimate of N Oph 2008 by W. Vollmann, Vienna, Austria: May 26.969 UT, 10.4. Photoelectric V magnitudes by L. Elenin, Lubertsy, Russia, contributed by E. Waagen, AAVSO: May 26.457, 10.30; 27.380, 10.14.

2008 May 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green